# Efficient Speech Translation through Model Compression and Knowledge Distillation

### v1

@@ 1 Introduction
Multimodal foundation models have shown powerful capabilities in different tasks, including speech translation. However, these models are often large and computationally intensive, making them impractical to use in real-world settings with limited resources. To enhance the efficiency and sustainability of these models, researchers have been investigating diverse approaches to model compression that aim to reduce the computational requirements of foundation models while retaining their performance (Treviso et al., 2023; Gandhi et al., 2023).
Qwen2-Audio (Chu et al., 2023, 2024) is a state-of-the-art foundation model that accepts various audio signal inputs and performs audio analysis or direct textual responses to speech instructions. In the IWSLT 2025’s Model Compression track, the organizes required that all submissions must be derived from the Qwen2-Audio model. The official languages of the task are English-to-German (EN-DE) and English-to-Chinese (EN-ZH).
@@
  • Model A: This model is the outcome of fine-tuning Qwen2-Audio-7B-Instruct in two stages: (a) full fine-tuning with the ACL 60/60 dataset, and (b) QLoRA fine-tuning with 4-bit quantization using the ACL 60/60 dataset augmented with data knowledge distillation from the fully fine-tuned model. This process has achieved 40% compression in terms of both model parameters and storage size. We discuss the details of this model in Section 3.2.
  • Model B: This model is a pruned version of Qwen2-Audio-7B-Instruct, and was created in multiple stages: (a) full fine-tuning of the baseline model with the ACL 60/60 dataset, (b) layer pruning of the decoder into 24 layers, while the encoder 32 layers were kept intact, (c) full fine-tuning of the pruned model, and (d) QLoRA fine-tuning with the ACL 60/60 dataset augmented with data knowledge distillation from the fully fine-tuned model and a portion of the CoVoST2 dataset to restore the quality of the teacher model. This process has achieved 50% compression in terms of both model parameters and storage size. We discuss the details of this model in Section 3.3.
@@
Table 1: Evaluation of the experiment where we use QLoRA with 4-bit Quantization, and augment the authentic data with knowledge distillation. This approach reduces the model size by more than 40% in terms of both the number of parameters (params) and storage footprint, while achieving the best translation performance for both English-to-German (EN-DE) and English-to-Chinese (EN-ZH) language pairs.
@@ 2.1 In-Domain Data
The ACL 60/60 dataset [^2] https://hf.co/datasets/ymoslem/acl-6060 (Salesky et al., 2023) is used in all of our experiments as the in-domain data. ACL 60/60 contains multilingual translation of ACL 2022 technical presentations into 10 target languages. The dataset consists of two splits, “dev” and “eval”, which together comprise 884 utterances. We merged the two splits, and randomly sampled 100 utterances for testing, which left us with 784 utterances for training. For test data sampling, we used the train_test_split method from the datasets [^3] https://github.com/huggingface/datasets library (Lhoest et al., 2021), setting the random seed option into zero. The ACL 60/60 dataset was required for “constrained” submissions to the IWSLT’s “Model Compression” track.
@@ 2.2 Out-of-Domain Data
After layer pruning of the Qwen2-Audio-7B-Instruct model (cf. Section 3), we needed to use more training data to restore the translation quality of the unpruned model. In addition to knowledge distillation data from the teacher model, we used a portion of the CoVoST2 dataset (Wang et al., 2021) (cf. Section 3.3). CoVoST2 is a large-scale multilingual speech-to-text translation corpus. Its latest version covers translations from English into 15 languages, including German and Chinese.
@@
Table 2: Evaluation of the layer pruning experiment. We started by full-parameter fine-tuning (Full FT) of the baseline model Qwen2-Audio-7B-Instruct on in-domain dataset ACL 60/60. Pruning 8 layers of the decoder of the model resulted in approx. 20% reduction in the model size; however, it affected the quality of the model. Hence, we fine-tuned the pruned model again on the in-domain dataset to restore as much as possible of the quality of the fully fine-tuned model. Finally, we fine-tuned the resulted model with low-rank adaptation after quantizing it into the 4-bit precision (QLoRA) on a mix of the in-domain data, knowledge distillation data (KD) and out-of-domain data, namely the CoVoST2 (CV) dataset. The QLoRA fine-tuning step resulted in approx. 50% reduction in the model size, while retaining 97% and 100% of the translation quality for Chinese and German, respectively, compared to the fully fine-tuned teacher model.
@@ 3 Methodologies
Our first experimental setup (cf. Section 3.2) employs QLoRA fine-tuning with 4-bit quantization, while the other experimental setup (cf. Section 3.3) conducts layer importance evaluation and applies iterative layer pruning of the model, followed by QLoRA fine-tuning. Both setups use knowledge distillation to retrieve the translation quality of the in-domain teacher model.
@@ 3.1 Full-Parameter Fine-tuning
In all of our experiments, we start by full-parameter fine-tuning of the Qwen2-Audio-7B-Instruct model on the ACL 60/60 dataset. This step is essential to make sure the foundation model is familiar with the downstream task and in-domain data. In particular, we train the baseline model on the in-domain data for 3 epochs, using a batch size of 4, learning rate of 1e-5, weight decay of 0.001, and no warm-up steps. The model is initially loaded in bfloat16 data type. As shown in Table 1 and Table 2, the fully fine-tuned model clearly outperforms the baseline by an average of 48 COMET points for translation into German and Chinese. Hence, this fully fine-tuned model is used as a foundation for both our experiments in Section 3.2 and Section 3.3. It is also used as a “teacher” for knowledge distillation into the compressed “student” models. We conduct the training on one H200 SXM GPU, using the Transformers framework [^4] https://github.com/huggingface/transformers (Wolf et al., 2020).
@@ 3.2 QLoRA with 4-bit Quantization
To restore the quality of the in-domain fully fine-tuned teacher models (cf. Section 3.1), sequence-level knowledge distillation is applied (Kim and Rush, 2016; Gandhi et al., 2023). In other words, we translate the ACL 60/60 training data with the “teacher” model. The training data is then augmented with the knowledge distillation data, and duplicates are filtered out. As a result, the augmented data after knowledge distillation comprises 1,568 for German and 1,069 for Chinese. Finally, “student” models are fine-tuned with QLoRA on the augmented data.
@@
As Table 1 illustrates, the combination of QLoRA with 4-bit quantization and knowledge distillation has achieved the highest translation performance into both German and Chinese across all evaluation metrics while reducing the model size in terms of both the number of parameters and storage requirements by more than 40% of the baseline model. [^5] For consistency, our storage footprint calculations are based only on the size of model files (*.safetensors), including the adapter of QLoRA models. We exclude the tokenizer and configuration files, as their contribution to the overall size is relatively small ($\sim$ 13 MB) and they are unaffected by our optimization methods. All sizes are computed using the decimal definition of gigabytes, where $1~{}\mathrm{GB}=1000^{3}$ bytes.
@@ 3.3 Iterative Layer Pruning
In this experimental setup, we apply iterative layer pruning to the fully fine-tuned “teacher” model (cf. Section 3.1). This approach incrementally identifies and removes layers with minimal contribution to translation quality, one layer at a time. The pruned model resulting from this process is then fine-tuned on both the ACL 60/60 training dataset and knowledge distillation data from the teacher model. Finally, the model is further fine-tuned with QLoRA, leveraging 4-bit quantization for efficient low-rank adaptation. Pruning 8 layers achieves a 20% reduction in model size, which increases to 50% when combined with 4-bit quantization. Fine-tuning the pruned model restores between 97% and 100% for German and Chinese, respectively, compared to the translation quality of the teacher model. The following points elaborate further on the process.
@@
This step serves two purposes, improving both the compression level and translation performance of the pruned model. After quantizing the model resulting from the previous step, low rank adaptation is used for fine-tuning it further. The training data consists of the ACL 60/60 dataset augmented with the knowledge distillation data from the fully fine-tuned “teacher” model (oversampled by a factor of 10), as well as a portion of the CoVoST2 dataset (100k utterances). We fine-tune the model for 1 epoch, and use a batch size of 8 (to make use of the computing resources, since the model is much smaller now), and a learning rate of 1e-5.
@@
Our ablation study (cf. Section 4.3) demonstrates that iterative layer pruning considerably outperforms fixed middle-layer pruning (cf. Section 4.3). Moreover, it clarifies that pruning exclusively decoder layers yields better performance than pruning both encoder and decoder layers (cf. Section 4.2). While our main experiments in this section prune only 8 layers, resulting in a model with 24 decoder layers and 32 encoder layers, the ablation study investigates pruning up to 16 layers (cf. Section 4.4).
Encoder Decoder BLEU chrF++ COMET Prams Storage
@@ 4.1 Qwen2-Audio Base vs Instruct
Language Data BLEU chrF/++ COMET
@@ 4.2 Encoder-Decoder Layer Pruning
The Qwen2-Audio is based on the encoder-decoder Transformer architecture (Vaswani et al., 2017). It consists of an encoder for audio (audio_tower) and a decoder for text generation (language_model), each of which comprises 32 layers. We first experimented with layer pruning of both the encoder and decoder. However, inspired by the Distil-Whisper work Gandhi et al. (2023), we experimented with pruning decoder layers only, which achieves better results. In other words, pruning only the decoder from 32 layers to 24 layers outperformed pruning both the encoder and decoder into 24 layers.
In this experiment, we pruned 8 fixed middle layers, from the 12th to the 19th layer, inclusively. After fine-tuning both models with the English-to-German ACL 60/60 dataset, the model where only the decoder layers were pruned achieved 4+ more points in terms of BLEU and chrF++ and 18+ points of COMET. It is worth noting that both the number of parameters and the storage footprint of this model is only 1% larger than the model with both the encoder and decoder are pruned. Table 3 shows the results of fine-tuning the English-to-German model after middle-layer pruning and fine-tuning.
@@
Table 5: Comparison between middle-layer pruning and iterative layer pruning, with either COMET or chrF/chrF++ as the metric for measuring layer importance. Iterative layer pruning, i.e. removing layers one by one, and then evaluating the resulted model, outperforms middle-layer pruning. In particular, when chrF/chrF++ is used for layer importance evaluation, the resulted pruned model achieves better speech translation quality after fine-tuning on the ACL 60/60 dataset.
@@ 4.3 Iterative vs Middle Layer Pruning
As discussed in Section 3.3, we experimented with iterative pruning based on layer importance evaluation to identify and remove the layers that contribute least to translation quality. On the contrary, in middle-layer pruning, we simply remove the 8 middle layers of the model, namely layers 12 to 19 inclusively out of the 32 decoder layers of the model.
@@
In iterative layer pruning based on performance evaluation (using chrF++ for German and chrF for Chinese), the pruned layers are [1, 3, 9, 13, 19, 20, 27, 29] for German, and [3, 4, 7, 15, 20, 25, 26, 29]. This shows a diverse layer selection that is not only concentrated in the middle. As Table 5 illustrates, iterative layer pruning yields much better results than middle-layer pruning. For example, when pruning 8 middle layers of the English-to-Chinese model, the final evaluation scores of the resulting model are so low across all metrics (BLEU: 1.3, chrF: 3.8, COMET -94.05), compared to the scores achieved when the same number of layers are iteratively removed based layer importance (BLEU: 42.52, chrF: 36.92, COMET 39.42).
Moreover, we experimented with both using chrF/chrF++ and COMET for evaluating pruned models during the iterative process. Interestingly, the final model resulting from using chrF/chrF++ for layer importance evaluation achieves better results. This might be due to the scientific nature of the ACL 60/60 dataset, and it requires future exploration with other datasets.
@@
Table 6: Comparison of iterative pruning of 8 decoder layers (which is the foundation of our pruning experiments) against pruning 10, 12, and 16 decoder layers. We observe that the quality of German degrades much more rapidly than that of Chinese, after pruning more than 8 layers. Moreover, pruning 16 layers from the Chinese model results in notable worse performance compared to pruning only 8, 10, or 16 layers, even after fine-tuning. When iteratively removing 16 layers from the German model, the removed layers are: [13, 3, 20, 9, 29, 1, 19, 27, 7, 26, 15, 18, 10, 17, 14, 30]. For the Chinese model, the removed layers are [7, 25, 3, 4, 29, 15, 26, 20, 24, 5, 10, 9, 27, 28, 18, 13]. The layers are listed from most to least important (left to right), according to the layer importance evaluation used in our iterative pruning method. In this experiment, all 32 encoder layers are kept intact; therefore, the table reports the remaining decoder layers as 24/32, 22/32, 20/32, and 16/32 encoder/decoder layers.

### v2

@@ 1 Introduction
Multimodal foundation models have shown powerful capabilities in different tasks, including speech translation. However, these models are often large and computationally intensive, making them impractical to use in real-world settings with limited resources. To enhance the efficiency of these models, researchers have been investigating diverse approaches to model compression that aim to reduce the computational requirements while retaining performance (Gandhi et al., 2023; Peng et al., 2023b, a; Treviso et al., 2023; Wang et al., 2023).
Qwen2-Audio (Chu et al., 2023, 2024) is a state-of-the-art foundation model that accepts various audio signal inputs and performs audio analysis or direct textual responses to speech instructions. In the IWSLT 2025’s Model Compression track (Abdulmumin et al., 2025), the organizers required that all submissions must be derived from the Qwen2-Audio model. The official languages of the task are English-to-German (EN-DE) and English-to-Chinese (EN-ZH).
@@
  • Setup 1: This model is the outcome of fine-tuning Qwen2-Audio-7B-Instruct in two stages: (a) full fine-tuning with the ACL 60/60 dataset, and (b) QLoRA fine-tuning with 4-bit quantization using the ACL 60/60 dataset augmented with data knowledge distillation from the fully fine-tuned model. This process has achieved 40% compression in terms of both model parameters and storage size. We discuss the details of this model in Section 3.2.
  • Setup 2: This model is a pruned version of Qwen2-Audio-7B-Instruct, and was created in multiple stages: (a) full fine-tuning of the baseline model with the ACL 60/60 dataset, (b) layer pruning of the decoder into 24 layers, while all 32 encoder layers were kept intact, (c) full fine-tuning of the pruned model, and (d) QLoRA fine-tuning with the ACL 60/60 dataset augmented with data knowledge distillation from the fully fine-tuned model and a portion of the CoVoST2 dataset to restore the quality of the teacher model. This process has achieved 50% compression in terms of both model parameters and storage size. We discuss the details of this model in Section 3.3.
@@
Table 1: Evaluation of the experiment that employs QLoRA with 4-bit quantization and augments the authentic data with knowledge distillation. This approach reduces the model size by more than 40% in terms of both the number of parameters (params) and storage footprint, while achieving the best translation performance for both English-to-German (EN-DE) and English-to-Chinese (EN-ZH) language pairs.
@@ 2.1 In-Domain Data
The ACL 60/60 dataset [^2] https://hf.co/datasets/ymoslem/acl-6060 (Salesky et al., 2023) is used in all of our experiments as the in-domain data. ACL 60/60 contains multilingual translation of ACL 2022 technical presentations into 10 target languages. The dataset consists of two splits, “dev” and “eval”, which together comprise 884 utterances. We merged the two splits, and randomly sampled 100 utterances for testing, which left us with 784 utterances for training. For test data sampling, we used the train_test_split method from the datasets [^3] https://github.com/huggingface/datasets library (Lhoest et al., 2021), setting the random seed option to zero. The ACL 60/60 dataset was required for “constrained” submissions to the IWSLT’s “Model Compression” track.
@@ 2.2 Out-of-Domain Data
After layer pruning of the Qwen2-Audio-7B-Instruct model (cf. Section 3), we needed to use more training data to restore the translation quality of the unpruned model. In addition to knowledge distillation data from the teacher model, we used a portion of the CoVoST2 dataset (Wang et al., 2021) (cf. Section 3.3). CoVoST2 is a large-scale multilingual speech-to-text translation corpus, covering translations from English into 15 languages, including German and Chinese.
@@
Table 2: Evaluation of the iterative layer pruning experiment. We started by full-parameter fine-tuning (Full FT) of the baseline model Qwen2-Audio-7B-Instruct on the in-domain dataset ACL 60/60. Pruning 8 layers of the decoder of the model achieved approx. 20% reduction in the model size; however, it affected the quality of the model. Hence, we fine-tuned the pruned model again on the in-domain dataset to restore as much as possible of the quality of the fully fine-tuned model. Finally, we fine-tuned the resulting model with low-rank adaptation after quantizing it into the 4-bit precision (QLoRA) on a mix of the in-domain data, knowledge distillation data (KD) and out-of-domain data, namely the CoVoST2 (CV) dataset. The whole process of pruning followed by QLoRA fine-tuning with 4-bit quantization has resulted in approx. 50% reduction in the model size, while retaining 97% and 100% of the translation quality for Chinese and German, respectively, compared to the teacher model.
@@ 3 Methodologies
Our first experimental setup (cf. Section 3.2) employs QLoRA fine-tuning with 4-bit quantization, while the other experimental setup (cf. Section 3.3) conducts layer importance evaluation and applies iterative layer pruning of the model, followed by QLoRA fine-tuning. Both setups use knowledge distillation to recover the translation quality of the in-domain teacher model.
@@ 3.1 Full-Parameter Fine-tuning
In all of our experiments, we start by full-parameter fine-tuning of the Qwen2-Audio-7B-Instruct model on the ACL 60/60 dataset. This step is essential to ensure the foundation model is familiar with the downstream task and in-domain data. In particular, we train the baseline model on the in-domain data for 3 epochs, using a batch size of 4, learning rate of 1e-5, weight decay of 0.001, and no warm-up steps. The model is initially loaded in bfloat16 data type. As shown in Table 1 and Table 2, the fully fine-tuned model clearly outperforms the baseline by an average of 48 COMET points for translation into German and Chinese. Hence, this fully fine-tuned model is used as a foundation for both our experiments in Section 3.2 and Section 3.3. It is also used as a “teacher” for knowledge distillation into the compressed “student” models. We conduct the training on one H200 SXM GPU, using the Transformers framework [^4] https://github.com/huggingface/transformers (Wolf et al., 2020).
@@ 3.2 QLoRA with 4-bit Quantization
To restore the quality of the in-domain fully fine-tuned teacher models (cf. Section 3.1), sequence-level knowledge distillation is applied (Kim and Rush, 2016; Gandhi et al., 2023). In other words, we translate the ACL 60/60 training data with the “teacher” model. The training data is then augmented with the knowledge distillation data, and duplicates are filtered out. As a result, the augmented data after knowledge distillation comprises 1,568 segments for German and 1,069 segments for Chinese. Finally, “student” models are fine-tuned with QLoRA on the augmented data.
@@
As Table 1 illustrates, the combination of QLoRA with 4-bit quantization and knowledge distillation has achieved the highest translation performance into both German and Chinese across all evaluation metrics while reducing the model size in terms of both the number of parameters and storage requirements by more than 40% compared to the baseline model. [^5] For consistency, our storage footprint calculations are based only on the size of model files (*.safetensors), including the adapter of QLoRA models. We exclude the tokenizer and configuration files, as their contribution to the overall size is relatively small ($\sim$ 13 MB) and they are unaffected by our optimization methods. All sizes are computed using the decimal definition of gigabytes, where $1~{}\mathrm{GB}=1000^{3}$ bytes.
@@ 3.3 Iterative Layer Pruning
In this experimental setup, we apply iterative layer pruning to the fully fine-tuned “teacher” model (cf. Section 3.1). This approach incrementally identifies and removes layers with minimal contribution to translation quality, one layer at a time. The pruned model resulting from this process is then fine-tuned on both the ACL 60/60 training dataset and knowledge distillation data from the teacher model. Finally, the model is further fine-tuned with QLoRA, leveraging 4-bit quantization for efficient low-rank adaptation. Pruning 8 layers achieves a 20% reduction in model size, which increases to 50% when combined with 4-bit quantization. Fine-tuning the pruned model restores between 97% and 100% of the teacher model’s translation quality for Chinese and German, respectively. The following points elaborate on the process.
@@
This step serves two purposes, improving both the compression level and translation performance of the pruned model. After quantizing the model resulting from the previous step, low rank adaptation is used for fine-tuning it further. The training data consists of the ACL 60/60 dataset augmented with the knowledge distillation data from the fully fine-tuned “teacher” model (oversampled by a factor of 10), as well as a portion of the CoVoST2 dataset (100k utterances). We fine-tune the model for 1 epoch, using a batch size of 8 (to make use of the computing resources, since the model is much smaller now), and a learning rate of 1e-5.
@@
Our ablation study (cf. Section 4) demonstrates that iterative layer pruning considerably outperforms fixed middle-layer pruning (cf. Section 4.3). Moreover, it clarifies that pruning exclusively decoder layers yields better performance than pruning both encoder and decoder layers (cf. Section 4.2). While our main experiments in this section prune only 8 layers, resulting in a model with 24 decoder layers and 32 encoder layers, the ablation study investigates pruning up to 16 layers (cf. Section 4.5).
Encoder Decoder BLEU chrF++ COMET Params Storage
@@ 4.1 Qwen2-Audio Base vs Instruct
Language Model BLEU chrF/++ COMET
@@ 4.2 Encoder-Decoder Layer Pruning
The Qwen2-Audio is based on the encoder-decoder Transformer architecture (Vaswani et al., 2017). It consists of an encoder for audio (audio_tower) and a decoder for text generation (language_model), each of which comprises 32 layers. We first experimented with layer pruning of both the encoder and decoder. However, inspired by the Distil-Whisper work Gandhi et al. (2023), we experimented with pruning decoder layers only, which achieved better results. In other words, pruning only the decoder from 32 layers to 24 layers outperformed pruning both the encoder and decoder into 24 layers.
In this experiment, we pruned 8 fixed middle layers, from the 12th to the 19th layer, inclusively. After fine-tuning both models with the English-to-German ACL 60/60 dataset, the model where only the decoder layers were pruned achieved 4+ additional points in terms of BLEU and chrF++ and 18+ points of COMET. It is worth noting that both the number of parameters and the storage footprint of this model is only 1% larger than the model with both the encoder and decoder were pruned. Table 3 shows the performance evaluation results of fine-tuning the English-to-German model after middle-layer pruning.
@@
Table 5: Comparison between middle-layer pruning and iterative layer pruning, with either COMET or chrF/chrF++ as the metric for measuring layer importance. Iterative layer pruning, i.e. removing layers one by one, and then evaluating the resulting model, outperforms middle-layer pruning. In particular, when chrF/chrF++ is used for layer importance evaluation, the resulting pruned model achieves better speech translation quality after fine-tuning on the ACL 60/60 dataset.
@@ 4.3 Iterative vs Middle Layer Pruning
As discussed in Section 3.3, we experimented with iterative pruning based on layer importance evaluation to identify and remove the layers that contribute least to translation quality. In contrast, in middle-layer pruning, we simply remove the 8 middle layers of the model, namely layers 12 through 19 out of the 32 decoder layers of the model.
@@
In iterative layer pruning based on performance evaluation (using chrF++ for German and chrF for Chinese), the pruned layers are [1, 3, 9, 13, 19, 20, 27, 29] for German, and [3, 4, 7, 15, 20, 25, 26, 29] for Chinese. This shows a diverse layer selection that is not concentrated solely in the middle. As Table 5 illustrates, iterative layer pruning yields much better results than middle-layer pruning. For example, when pruning 8 middle layers of the English-to-Chinese model, the final evaluation scores of the resulting model are so low across all metrics (BLEU: 1.3, chrF: 3.8, COMET -94.05), compared to the scores achieved when the same number of layers is iteratively removed based on layer importance (BLEU: 42.52, chrF: 36.92, COMET 39.42).
Moreover, we experimented with both using chrF/chrF++ and COMET for evaluating pruned models during the iterative process. Interestingly, the final model obtained by using chrF/chrF++ for layer importance evaluation achieves better results (cf. Table 5). This might be due to the scientific nature of the ACL 60/60 dataset, and it requires future exploration with other datasets.
@@
Table 6: Comparison of iterative pruning of 8 decoder layers (which is the foundation of our pruning experiments) against pruning 10, 12, and 16 decoder layers. We observe that the quality of German degrades much more rapidly than that of Chinese, after pruning more than 8 layers. Moreover, pruning 16 layers from the Chinese model results in notably worse performance compared to pruning only 8, 10, or 12 layers, even after fine-tuning. When iteratively removing 16 layers from the German model, the removed layers are: [13, 3, 20, 9, 29, 1, 19, 27, 7, 26, 15, 18, 10, 17, 14, 30]. For the Chinese model, the removed layers are [7, 25, 3, 4, 29, 15, 26, 20, 24, 5, 10, 9, 27, 28, 18, 13]. The layers are listed from most to least important (left to right), according to the layer importance evaluation used in our iterative pruning method. In this experiment, all 32 encoder layers are kept intact; therefore, the table reports the remaining decoder layers as 24/32, 22/32, 20/32, and 16/32 encoder/decoder layers.
4.4 Immediate Recovery
In our main experiments, we fine-tuned the pruned models only after completing the entire pruning process (cf. Section 3.3). In order to understand the effect of accompanying iterative pruning with iterative recovery (Wibowo et al., 2025), we conducted an extra experiment where we immediately fine-tuned the model after each layer pruning iteration. In other words, when pruning 8 layers, the fine-tuning is performed 8 times, once after each pruning step.
By the end of the process, the pruned layers for German are [13, 1, 3, 17, 12, 4, 14, 21], based on layer importance evaluation. These layers differ from those selected when pruning without immediate recovery (cf. Section 4.3), since fine-tuning after each iteration changes the relative importance of the remaining layers.
Model BLEU chrF++ COMET
pruned 10.78 39.58 -44.40
+ FT [after] 32.16 60.39 39.08
pruned 30.96 59.02 30.07
[…]
While this immediate recovery approach improved evaluation scores of the pruned model, it did not achieve performance gains over our standard method of fine-tuning only once after complete pruning. This might be due to overfitting caused by fine-tuning several times on the small in-domain dataset. Moreover, immediate recovery through fine-tuning after each pruning iteration is much more computationally intensive.